OpenStudy (anonymous):

factor completely x^2-9x+8

OpenStudy (zzr0ck3r):

we want two numbers that multiply to get 8 but add to get -9, what are these numbers?

OpenStudy (anonymous):

not sure

OpenStudy (anonymous):

-8 -1

OpenStudy (zzr0ck3r):

yes

OpenStudy (anonymous):

does it matter which i put in which

OpenStudy (zzr0ck3r):

ok now expand the middle term x^2-9x+8 = x^2-x-8x+8 factor out what we can from the first two terms x(x-1)-8x+8 now what can we factor from the 2nd two terms to make it x-1? factor out -8 x(x-1)-8(x-1) factor out x-1 (x-1)(x-8)=x^2-9x+8

OpenStudy (anonymous):

I already gpt the answer

OpenStudy (zzr0ck3r):

once you had the two numbers, because we have no coefficient on the x^2 term we could have just said(x-1)(x-8)

OpenStudy (zzr0ck3r):

but if you follow what I did, you can do any of these...

OpenStudy (anonymous):

what about with a more complicated one, 6x^2 - 7xy -3x^2

OpenStudy (anonymous):

x^2-8x-x+8 x(x-8)-1(x-8) (x-1)(x-8) x=1,x=8

OpenStudy (zzr0ck3r):

what is 6x^2-3x^2?

OpenStudy (anonymous):

3x^2

OpenStudy (zzr0ck3r):

so you have 3x^2-7xy do either of the two terms share anything?

OpenStudy (anonymous):

y?

OpenStudy (anonymous):

3x^2-7xy x(3x-7y)

OpenStudy (zzr0ck3r):

are there any y's in the first term 3x^2?

OpenStudy (zzr0ck3r):

@farhan2h we don't need two people doing the same thing, I am trying to lead her to the answer...not just paste it.

OpenStudy (anonymous):

i meant x sorry

OpenStudy (zzr0ck3r):

ok so factor the x out and what do you have?

OpenStudy (anonymous):

what the other guy said

OpenStudy (zzr0ck3r):

yep

OpenStudy (anonymous):

okay I get it

OpenStudy (zzr0ck3r):

ok but notice that this is way easier than the other question, and if this had some other terms, it could have gotten very very ugly very very fast.

OpenStudy (anonymous):

i have one with other terms

OpenStudy (anonymous):

2u^2 + uv - 21v^2

OpenStudy (zzr0ck3r):

ok so this is more like the first one, we want two numbers that multiply to -42 and add to 1(the middle term has a 1 in front of it) two such numbers are -6,7 so expand the middle term 2u^2+uv-21v^2 = 2u^2-6uv+7uv-21v^2 what can we factor out of the first two terms?

OpenStudy (anonymous):

u

OpenStudy (zzr0ck3r):

2u

OpenStudy (anonymous):

okok

OpenStudy (zzr0ck3r):

then we have 2u(u-3v)+7uv-21v^2 what can we factor out of the last two terms?

OpenStudy (anonymous):

-21 v

OpenStudy (zzr0ck3r):

not 21 but 7

OpenStudy (zzr0ck3r):

you see why?

OpenStudy (anonymous):

no

OpenStudy (zzr0ck3r):

2u(u-3v)+7uv-21v^2 factor out 7u because each of the last two terms can be divided by 7u

OpenStudy (zzr0ck3r):

if you had (21x+7) you can factor out a 7 not a 21 7(3x+1)

OpenStudy (anonymous):

oh okay

OpenStudy (zzr0ck3r):

I mean 7v

OpenStudy (zzr0ck3r):

sorry

OpenStudy (zzr0ck3r):

and we want it so that what we have left after we factor the last two terms looks like the thing in the first two terms 2u(u-3v)+7uv-21v^2 factor out 7v because each of the last two terms can be divided by 7v 2u(u-3v)+7v(u-3v) let me know when you understand up to here.

OpenStudy (anonymous):

i got it

OpenStudy (zzr0ck3r):

ok notice now we have u-3v in both terms?

OpenStudy (zzr0ck3r):

factor that out 2u(u-3v)+7v(u-3v) = (u-3v)(2u+7v)

OpenStudy (zzr0ck3r):

this is the same thing as 2ux+7vx = x(2u+7v) but instead of x we have u-3v so 2u(u-3v) + 7v(u-3v) = (u-3v)(2u+7v)

OpenStudy (zzr0ck3r):

what we did 1) find two numbers that multiple to our coefficient on the last term, but add to the coefficient on the middle term 2) expand that middle term with out new numbers 3) factor what we can out of the first 2 terms 4) factor what we can out of the last 2 terms such that we are left with the same form as one of the factors from part 3) 5) then factor that thing out I know this is not the best mathematical terms, but if you can do this with all the quadratics you get(that are factorable), and if you understand this your good...